y-x=3x
2(x+y)=14-y

3a+6b=-15
a=-2b-5
pls. solve by substitution

Answer by mananth(16946)

y-x=3x
add x to both sides
y= 4x

2(x+y)=14-y
substitute y
2(x+4x)=14+3x
10x=14+3x
add -3x to both sides
7x=14
/7
x=2
when x =2 y = 4x = 4*2=8
(2,8)